Cans of tomato sauce are to be cooled in a refrigerator. The cans are cylindrical and 20 cm in length and 15 cm in diameter. The can surface is at 30°C and the air in the refrigerated space is at 10°C. Two alternate stacking are possible-namely vertical stacking and horizontal stacking. Determine the method that will be better. Neglect cooling on the end surfaces


1
Expert's answer
2021-12-14T09:27:48-0500

Horizontal stacking is better,

"S_h=\pi dh=940~cm^2,"

"S_v=\pi d^2=707~cm^2,"

"S_h>S_v."